Assume your U.S. company has a contract to purchase 100,000 computers from a Korean company. The payment is due on receipt of the shipment and must be delivered in Korea on December 31, 2018. In July 2018, when you are arranging the contract, the computers are priced at 500,000 won each. The spot rate in July 2018 is $1 in exchange for 1,088 won.

1. Calculate the U.S. dollar price (in July 2018) of one unit of Korean currency.

2. What is the total price of the computers in dollars?
3. What is the total price of the computers in won?
Answer #1

a)

Cost of one unit of laptop = 7000 won

Exchange rate = $1 in exchange for 18.77 won

USD price of one unit of laptop = 7000/18.77 = $372.94

b)

If won appreciates by 10%

New exchange rate = $1 in exchange for 16.89 won

USD price of one unit of laptop = 7000/16.89 = $414.45

Increase in USD laptop price = 414.45 - 372.94 = $41.51

Loss on purchase of 10000 computers = 10000*$41.51 = $415100